OpenStudy (anonymous):

what is the derivative of: 4.15^(cos(x))

OpenStudy (anonymous):

Use the chain rule: \[f(x) = 4.15^{\cos(x)}\] \[f(x)\prime = (4.15^{\cos(x)}*\ln(4.15)) * -\sin(x)\]